# Online Slip Detection and Friction Coefficient Estimation for Autonomous Racing

### v2

@@ V-C Ground-Truth Friction Measurement
To determine ground-truth values for the static friction coefficient on each surface, we conducted 42 independent trials, using a force gauge at the vehicle’s center of mass to pull the vehicle until it slipped. Seven repeated pulls were applied in each of three directions: lateral, longitudinal, and diagonal ($45^{\circ}$). This process was repeated for each surface.
Tire-road friction coefficients were then calculated based on the maximum force observed before slipping:
$\mu^{\ast}=\frac{F_{\text{pull}}}{F_{\text{normal}}}$ (20)
The ground-truth results are shown in gray in Figure 4. The experiments showed a slight directional dependence in the friction force. Cardboard maintains high friction across orientations ($\mu\approx 1.02$$1.04$), whereas tile exhibits lower values ($\mu\approx 0.66$$0.72$), consistent with expectations and demonstrating minor anisotropy due to surface texture.
V-D Question 1: Slip Detection Accuracy
@@
Given that we have labels for the times that the vehicle has slipped, we evaluated the accuracy of our slip detector. The slip detector performed well across our dataset, with a precision of 0.909, a recall of 1.000, and an F1-score of 0.952. Our approach detected every drift in our dataset, with no false negatives but two false positives that contributed to the lower precision.
V-E Question 2: Slip Detection Delay
To evaluate the delay of our slip detection module, we compared the times from our labeled bags with the times our approach detected. Our approach had a low absolute difference between the time of drift and the time of detection. The average absolute difference of tile surface detections was $0.243\pm 0.302$ seconds, and the average absolute difference of cardboard surface detections was $0.223\pm 0.112$ seconds.
V-F Question 3: Friction Estimation Accuracy
To determine the accuracy of our friction coefficient estimation, we compared the measured ground truth TRFC values with those from our approach. Our estimated values were $\mu\approx 0.757$ on tile and $\mu\approx 1.086$ on cardboard. These values were very accurate compared to the ground truth values, with a mean average error (MAE) of $0.089\pm 0.055$ for the tile surface and $0.157\pm 0.172$ for the cardboard surface (visualized in blue Figure 4).
Figure 5 shows an example of our TRFC estimation results on cardboard and tile surfaces. The red circle represents the friction circle, indicating the limit of possible tire force combinations. The blue dots indicate the measured longitudinal and lateral accelerations (in terms of g) during testing.
@@ V-F Question 3: Friction Estimation Accuracy
Figure 4: Comparison of ground truth (gray) and estimated (blue) friction coefficients ($\mu$) on tile and cardboard.
(a) Tile, $\mu\approx 0.844$
(b) Cardboard, $\mu\approx 1.209$
Figure 5: Depiction of the friction circle for each surface in one test bag. The blue dots represent the IMU data, while the star represents the maximum point, corresponding to the friction circle boundary (and thus $\mu$).
@@ VI DISCUSSION AND CONCLUSION
Nevertheless, several limitations must be acknowledged. The experiments were restricted to two surfaces and a limited number of trials, which constrains the generalizability of the results. Environmental factors such as humidity, wear of the surfaces, and variation in the manual pulling force applied to the vehicle’s center of mass during ground-truth measurements were not fully controlled, potentially influencing repeatability. Additionally, the current method assumes largely isotropic surfaces and enforces simplifying assumptions introduced in Section II — including planar dynamics, negligible aerodynamic drag and rolling resistance, and single-track approximation. These assumptions may not hold for textured or anisotropic terrains encountered in real-world deployments.
There are several promising directions for future work. Expanding the range of tested surfaces, especially those with irregular or compliant properties, would strengthen robustness. Relaxing the simplifying assumptions — for example, accounting for aerodynamic drag at higher speeds, modeling rolling resistance, or incorporating pitch/roll dynamics — would improve fidelity and extend applicability to more diverse conditions. Integration of real-time adaptive control using friction estimates could enhance navigation in varied-friction environments. Furthermore, coupling the presented techniques with learning-based models may allow for adaptive friction prediction under varying conditions. Finally, scaling the experiments to larger sample sizes and different platforms will be necessary to validate broader applicability.
In summary, this study provided an initial validation of onboard friction estimation techniques, showing strong alignment with external ground-truth methods. While the scope was intentionally narrow, the findings demonstrate that even lightweight onboard approaches can yield reliable surface characterizations, laying the groundwork for more comprehensive future investigations.

### v3

@@ V-C Ground-Truth Friction Coefficient Measurement
To determine ground-truth TRFC values on each surface, we conducted a total of 45 independent trials using a force gauge at the vehicle’s center of mass to pull the vehicle until it slipped. For each of the three surfaces, five repeated pulls were applied in each of three directions: lateral, longitudinal, and diagonal ($45^{\circ}$). For the diagonal direction, pulls were applied from the front of the vehicle while alternating between the two diagonal orientations ($45^{\circ}$ and $45^{\circ}$) to ensure balanced coverage.
Measured static friction values were approximately $\mu\approx 0.69\pm 0.03$ (tile), $\mu\approx 1.02\pm 0.02$ (cardboard), and $\mu\approx 0.84\pm 0.03$ (acrylic) as shown in gray in Figure 5.
V-D Results
Question 1: Slip Detection Accuracy
@@ V-D Results
We evaluated the accuracy of our slip detector by comparing its detected slip events against labeled ground-truth slip times. The slip detector performed exceptionally well across our dataset, with a precision of 0.987, a recall of 1.000, and an F1-score of 0.993. That is, our approach detected every drift in our dataset (148 events), with no false negatives and only two false positives. Both occurred on track 1 on ceramic tile, near the end of the corresponding bags, where a light contact with the wall contributed to the spurious detections.
Question 2: Slip Detection Delay
We measured the delay of our slip detection approach by comparing the timestamp of each detection event with the visual onset of slipping recorded in the video of each lap. Our approach demonstrated a small difference between the time of drift and the time of detection. The average absolute difference of tile surface detections was $0.407\pm 0.534$ s, the average absolute difference of cardboard surface detections was $0.476\pm 0.496$ s, and finally the average absolute difference of the acrylic was $0.579\pm 0.609$ s.
@@
Figure 5: Comparison of ground truth (gray) and estimated (blue) friction coefficients (µ) on tile, cardboard, and acrylic.
Question 3: Friction Estimation Accuracy
To determine the accuracy of our friction coefficient estimation, we compared the measured ground-truth TRFC values with those produced by our approach. Our estimated values were $\mu\approx 0.714$ on tile, $\mu\approx 1.038$ on cardboard, and $\mu\approx 0.860$ on acrylic. These values were, on average, accurate compared to the ground truth values, with a mean absolute error (MAE) of $0.042\pm 0.032$ for the tile surface, $0.116\pm 0.120$ for the cardboard surface, and $0.139\pm 0.097$ for the acrylic surface (visualized in blue in Figure 5). As expected, our approach showed some variance due to sensor noise. Figure 4 shows an example of our TRFC estimation results on cardboard, acrylic, and tile surfaces. The red circle represents the friction circle, indicating the limit of possible tire force combinations where maximum friction is estimated. The variability of acceleration demonstrates that our detection and estimation approach was tested comprehensively.
Table I contextualizes our experimental results in terms of existing TRFC estimation approaches. The metrics include MAE and the mean relative error (MRE, i.e., MAE normalized by the true TRFC). Note that for the adaptive cubature KF [39], MAE values were not explicitly reported; therefore, we approximated them by digitizing the published plots and extracting the corresponding error values. Recall that, unlike most model-based approaches, which rely on explicit vehicle or tire models and parameter identification, and unlike learning-based methods that require extensive training data, our approach operates without parametric modeling or prior training, using only IMU and LiDAR measurements together with control inputs. Despite this lightweight formulation, the achieved estimation errors remain within a comparable range to filtering-based techniques. In particular, across multiple surface conditions, our method maintains consistent performance without surface-specific tuning or calibration. These results support the central claim of this work: accurate friction estimation can be achieved through a simple, model-free, and computationally efficient framework suitable for embedded onboard deployment.
TABLE I: Error comparison for tire–road friction coefficient estimation methods. MRE denotes MAE normalized by the ground-truth.
Method (paper) MAE MRE (%)
Deep-learning based [32] 0.019 2.77
Refined adaptive KF [19] 0.142 17.75
0.02 3.6
0.01 2.3
Predictive ELSD Control (multiple surfaces) [40] 0.02 8.3
0.01 5.5
0.02 2.2
Adaptive cubature KF (approximation) [39] 0.127 15.9
Wavelet transform and neural network [23] 0.043 5.30
0.042 6.09
Ours (multiple surfaces) 0.116 11.37
0.139 16.45
@@ VI DISCUSSION AND CONCLUSION
Nevertheless, several limitations must be acknowledged. The experiments were restricted to three surfaces and a limited number of trials, which constrains the generalizability of the results. Environmental factors such as humidity, wear of the surfaces, and variation in the manual pulling force applied to the vehicle’s center of mass during ground-truth measurements were not fully controlled, potentially influencing repeatability. Additionally, the current method assumes largely isotropic surfaces and enforces simplifying assumptions introduced in Section II — including planar dynamics, negligible aerodynamic drag and rolling resistance, and single-track approximation. These assumptions may not hold for significantly textured or anisotropic terrains encountered in real-world deployments.
Future work will focus on extending validation beyond controlled indoor surfaces to more diverse and dynamically changing conditions, such as outdoor asphalt and wet or low-adhesion environments. Incorporating additional vehicle dynamics effects, including aerodynamic drag and load transfer, would further improve estimation accuracy during higher-speed or aggressive maneuvers. Another direction is improving slip onset localization using sensor-driven criteria (encoder-based) rather than visual labeling, enabling more precise delay analysis. Finally, hybrid extensions that combine the present physics-based framework with lightweight learning components may enhance generalization across unseen surfaces while preserving the interpretability and computational efficiency of the current approach.
This study provided an initial validation of onboard friction estimation techniques, showing strong alignment with external ground-truth methods. While the scope was intentionally narrow, the findings demonstrate that even lightweight onboard approaches can yield reliable surface characterizations, laying the groundwork for more comprehensive future investigations.